# A Survey on Large Language Model based Human-Agent Systems

### v1

@@ 7 Conclusion
This paper presents a comprehensive review of LLM-based Human-Agent Systems. We introduce a structured taxonomy covering five core dimensions, environment and profiling, human feedback, interaction types, orchestration paradigms, and communication, and use it to classify and analyze existing research on LLM-HAS. We also summarize representative implementation frameworks, benchmark datasets, and evaluation schemes to support reproducibility and comparative analysis. Finally, we identify key challenges and unresolved issues in current LLM-HAS research. These issues remain major obstacles to the development of effective, adaptive, safe and trustworthy human-agent systems. We hope this review offers a comprehensive understanding of the LLM-HAS landscape and serves as a practical guide for future research.
@@ Appendix B Human Feedback Type and Subtype
In this appendix, we present a detailed introduction of human feedback types and their subtypes as shown in Table B.1. Table B.1 summarizes its corresponding definition and explains how human feedback guides or constrains an LLM-based agent’s learning process. While the main text has already discussed the broad categories of evaluative, corrective, guidance, and implicit of the human feedback to the LLM-based agent in interaction, here we unpack each category into more fine-grained forms, ranging from scalar ratings and preference rankings to direct edits, demonstrations, and inferred behavioral signals. The subtype can help us understand how we react with the LLM-agent with clear instruction and task definition. By acquiring this knowledge, the human are able to improve the quality of interaction with the LLM-based agent. In addition, this comprehensive breakdown enables a systematic comparison across current studies and highlights the diverse ways in which human users can steer, correct, or collaborate with the LLM-based-agent.
@@
Human Action/Control Human directly takes actions and control. Collaborate with humans to effectively finish tasks or learns from human actions.
Table B.1: Human Feedback Type and Subtype.The subtype of evaluative feedback includes preference ranking, scalar rating, and binary assessment. The subtype of corrective feedback includes the direct edits or refinement. The subtype of guidance feedback includes the demonstration and instructions or critiques. The subtype of implicit feedback include the human action or control.
@@ Appendix C Evaluation Metrics
Evaluating LLM-based human agent systems requires comprehensive methodologies that capture both objective performance metrics and subjective user experiences. Current evaluation strategies have evolved to address these multidimensional aspects, adopting diverse approaches tailored to different contexts and system designs. This section discusses these evaluation methods under three primary categories: Quantitative Evaluation, Qualitative Evaluation, and Mixed-Method Evaluation.
C.1 Quantitative Evaluation
Quantitative evaluations focus on objective metrics to systematically assess system performance across various tasks and frameworks. For instance, in the healthcare domain,  Vollmuth et al. (2023),  Van Leeuwen et al. (2022) leverage precision, recall, and F1-score metrics to evaluate AI-assisted diagnostic tasks, specifically in oncology and radiology. Similarly, financial domains employ quantitative metrics such as true positives and false positives to for fraud detection evaluation as exemplified in Al-Fatlawi et al. (2024). In manufacturing, quantitative performance metrics are emphasized for assessing AI assistance in process efficiency and safety compliance, as demonstrated by  Sankaran et al. (2022) and  Massaro (2022). Other specialized domains also propose tailored quantitative metrics, such as specific performance scores in game-based AI  Siu et al. (2021) and evaluation frameworks in neuromorphic computing  Yik et al. (2025).
C.2 Qualitative Evaluation
Complementing quantitative metrics, qualitative evaluations aim to examine subjective aspects such as user perceptions, trust, adaptability, and ethical considerations. For example,  Timmons et al. (2023) employs qualitative methods including interviews and case studies to investigate potential biases in mental health AI applications.  Rezwana and Maher (2023) explores qualitative feedback to assess the impact of AI on creative workflows, underscoring the importance of human-AI interaction dynamics.  Sharma et al. (2023) demonstrates the value of qualitative insights by examining conversational empathy improvements in AI-assisted mental health platforms through user feedback and thematic analyses. These qualitative methodologies are vital for uncovering the nuanced human factors influencing system adoption and effectiveness.
C.3 Mixed-Method Evaluation
The mixed-method approach integrates quantitative and qualitative evaluations, offering a holistic assessment tailored to specific contexts. This approach addresses the limitations inherent in exclusively quantitative or qualitative methods by combining measurable performance outcomes with rich user-centric insights. For example, Arias-Rosales (2022) evaluates AI-generated design output by pairing quantitative shape metrics with qualitative user assessments, providing deeper insights into subjective aesthetic values. Similarly, in finance,  Chakravorti et al. (2022) combines detection accuracy with user trust evaluations to identify issues related to transparency and interpretability. Mixed-method evaluations are particularly effective in understanding complex interactions between humans and AI, facilitating nuanced and contextualized evaluations that are essential across diverse domains. By leveraging the advantage of the mixed-method,  Fragiadakis et al. (2024) provides a robust and comprehensive framework that integrates multiple evaluation dimensions to effectively assess Large Language Model-based Human-Agent Systems across diverse contexts.
@@ Appendix D Tables
Table D.1: The ① Humans-Agent Configuration ② Human Feedback in LLM-based human–agent systems.
@@
Collaborative Gym (Shao et al., 2024) Arxiv’25 Link Single Single Corrective Refinement Segment During Task
@@
MTOM (Zhang et al., 2024b) Arxiv’24 Single Single Corrective Refinement Segment During Task
@@
FineArena (Xu et al., 2025) Arxiv’25 Single Multiple Guidance Demonstration Segment During Task
@@
Prison Dilemm (Jiang et al., 2025) Arxiv’25 Single Single Implicit User Action Segment During Task
@@
InteractGen (Sun et al., 2024b) THU’24 Multiple Multiple Guidance Demonstration Segment During Task
AI Chains (Wu et al., 2022) CHI’24 Single Single Corrective Refinement Segment During Task
Drive As You Speack (Cui et al., 2024) WACV’24 Single Multiple Guidance Demonstration Segment During Task
@@
AgentCoord (Pan et al., 2024a) Arxiv’24 Link Single Multiple Guidance, Corrective Demonstration, Refinement Segment Initial Setup
@@
CowPilot (Huq et al., 2025) Arxiv’25 Link Single Single Corrective, Implicit User Action, Refinement Segment During Task
@@
EasyLAN (Pan et al., 2024c) Arxiv’24 Single Multiple Corrective, Guidance Demonstration, Refinement Segment During Task
@@
Hierarchical Agent (Liu et al., 2023b) AAMAS’24 Single Multiple Guidance, Corrective Demonstration, Refinement Segment During Task
@@
SWEET-RL (Zhou et al., 2025) Arxiv’25 Link Single Single Corrective, Implicit Refinement, User Action Segment During Task
@@
HRC Assembly (Gkournelos et al., 2024) CIRP’24 Single Multiple Guidance Demonstration Segment During Task
REVECA (Seo et al., 2025) Arxiv’24 Single Single Implicit Human Control Segment During Task
AssistantX (Sun et al., 2024a) Arxiv’24 Link Multiple Multiple Implicit Human Control Holistic Initial Setup
MINT (Wang et al., 2024c) ICLR’24 Link Single Single Evaluative Binary Assessment Holistic During Task
Help Feedback (Mehta et al., 2024) EACL’24 Single Single Corrective, Guidance Demonstration, Refinement Holistic During Task
ConvCodeWorld (Han et al., 2025) ICLR’25 Link Single Single Guidance Demonstration, Critique Segment During Task
ReHAC (Feng et al., 2024) ACL’24 Link Single Single Implicit Human Control Segment During Task
DPT Agent (Zhang et al., 2025) Arxiv’25 Link Single Single Guidance Critique Holistic During Task
HRC Manipulation (Liu et al., 2023a) IEEE’23 Single Single Corrective, Guidance Demonstration, Refinement Segment During Task
HRC DMP (Liu et al., 2024) IEEE’24 Single Single Corrective Refinement Holistic During Task
PARTNR (Chang et al., 2024) ICLR’25 Link Single Single Corrective, Guidance Refinement, Critique Holistic, Segment During Task, Post Task
Organized Teams (Guo et al., 2024b) Arxiv’24 Link Single Multiple Guidance Critique Holistic During Task
CoELA (Zhang et al., 2024a) ICLR’23 Single Multiple Evaluative Scaler Rating Holistic Post Task
Agency Task (Sharma et al., 2024) EACL’24 Link Single Single Guidance Demonstration, Critique Segment During Task
GDfC (Wang et al., 2025c) SME’25 Single Multiple Guidance, Evaluative Demonstration, Binary Assessment, Preference Ranking Holistic, Segment Initial Setup, Post Task
PDFChatAnnotator (Tang et al., 2024) IUI’24 Single Single Corrective, Guidance Demonstration, Refinement Segment During Task
Attentive Supp. (Tanneberg et al., 2024a) IEEE’24 Link Multiple Single Implicit, Guidance Demonstration, User Action Holistic Initial Setup, During Task
HRC Trust (Ye et al., 2023) IEEE’23 Single Single Guidance Demonstration, Critique Segment During Task
@@
HA Comm. (Bansal et al., 2024) Arxiv’24 Single Multiple Guidance Demonstration, Critique Holistic, Segment Initial Setup, During Task, Post Task
@@
BPMN (Ait et al., 2024) Arxiv’24 Link Multiple Multiple Guidance Demonstration Holistic Initial Setup, During Task, Post Task
Co-STORM (Jiang et al., 2024) EMNLP’24 Link Single Multiple Guidance Demonstration Segment During Task
HRC Manufa. (Lim et al., 2024) IEEE’24 Single Single Corrective, Guidance Demonstration, Refinement, Critique Segment Initial Setup, During Task
A2C (Tariq et al., 2025) Arxiv’24 Link Multiple Multiple Guidance, Implicit, Corrective, Evaluative Refinement, Binary Assessment, Critique, Human Control Holistic, Segment During Task
MindAgent (Gong et al., 2023) NAACL’24 Link Multiple Multiple Corrective Refinement Segment During Task
Ask Before Plan (Zhang et al., 2024c) EMNLP’24 Link Single Multiple Guidance Demonstration, Critique Segment During Task
SOTOPIA (Zhou et al., 2024) ICLR’24 Multiple Multiple Evaluative, Implicit Scaler Rating, User Action Holistic, Segment During Task, Post Task
PaLM-(Driess et al., 2023) ICML’23 Link Single Single Guidance, Implicit Demonstration, User Action Segment During Task
TaPA (Wu et al., 2023) Arxiv’23 Link Single Single Guidance, Evaluative Demonstration, Binary Assessment Holistic, Segment Initial Setup, Post Task
@@
MetaGPT (Hong et al., 2023) ICLR’24 Link Single Multiple Evaluative, Guidance Binary Assessment Holistic Initial Setup, Post Task
@@
DigiRL (Bai et al., 2024) NeurIPS’24 Link Single Single Evaluative Binary Assessment Holistic During Task, Post Task
@@
WebLINX (Lù et al., 2024) Arxiv’24 Link Multiple Single Guidance Demonstration Holistic, Segment During Task
@@
Auto Agent (Pan et al., 2024b) COLM’24 Link Single Single Evaluative Binary Assessment Holistic Post Task
WebCanvas (Pan et al., 2024d) Arxiv’24 Link Single Single Evaluative Scaler Rating Holistic Post Task
MineWorld (Guo et al., 2025) Arxiv’25 Link Multiple Single Evaluative Scaler Rating Holistic Post Task
Table D.2: The ① Interaction ② Orchestration ③ Communication in LLM-based human–agent systems.
@@
Paper Venue Code/ Data Types Variant Strategy Sync Structure Mode
Collaborative Gym (Shao et al., 2024) Arxiv’25 Link Collaboration Supervision, Delegation One-by-One Asynchronous Decentralized Conversation
MTOM (Zhang et al., 2024b) Arxiv’24 Collaboration Coordination Simultaneous Synchronous Decentralized Conversation
FineArena (Xu et al., 2025) Arxiv’25 Collaboration Delegation One-by-One Synchronous Hierarchical Conversation
Prison Dilemm (Jiang et al., 2025) Arxiv’25 Coopetition One-by-One Synchronous Decentralized Conversation
@@
AI Chains (Wu et al., 2022) CHI’24 Collaboration Delegation One-by-One Synchronous Hierarchical Conversation
Drive As You Speack (Cui et al., 2024) WACV’24 Collaboration Delegation One-by-One Synchronous Centralized Conversation
AgentCoord (Pan et al., 2024a) Arxiv’24 Link Collaboration Coordination One-by-One Synchronous Decentralized Conversation
@@
EasyLAN (Pan et al., 2024c) Arxiv’24 Collaboration Delegation, Supervision One-by-One Synchronous Hierarchical Observation
Hierarchical Agent (Liu et al., 2023b) AAMAS’24 Collaboration Supervision, Delegation Simultaneous Synchronous Hierarchical Conversation
SWEET-RL (Zhou et al., 2025) Arxiv’25 Link Collaboration Delegation One-by-One Synchronous Decentralized Conversation
HRC Assembly (Gkournelos et al., 2024) CIRP’24 Collaboration Delegation One-by-One Synchronous Centralized Conversation
REVECA (Seo et al., 2025) Arxiv’24 Collaboration Delegation One-by-One Asynchronous Hierarchical Observation
AssistantX (Sun et al., 2024a) Arxiv’24 Link Collaboration Delegation One-by-One Asynchronous Decentralized Message Pool
MINT (Wang et al., 2024c) ICLR’24 Link Collaboration Delegation One-by-One Synchronous Decentralized Conversation
Help Feedback (Mehta et al., 2024) EACL’24 Collaboration Delegation One-by-One Synchronous Decentralized Conversation
@@
ReHAC (Feng et al., 2024) ACL’24 Link Collaboration Delegation One-by-One Synchronous Decentralized Conversation
DPT Agent (Zhang et al., 2025) Arxiv’25 Link Collaboration Simultaneous Asynchronous Decentralized Observation
@@
HRC DMP (Liu et al., 2024) IEEE’24 Collaboration Delegation One-by-One Synchronous Decentralized Conversation
PARTNR (Chang et al., 2024) ICLR’25 Link Collaboration Coordination, Cooperation One-by-One Asynchronous Decentralized Conversation
Organized Teams (Guo et al., 2024b) Arxiv’24 Link Collaboration Delegation One-by-One Synchronous Decentralized Conversation
CoELA (Zhang et al., 2024a) ICLR’23 Collaboration Cooperation Simultaneous Synchronous Decentralized Conversation
@@
HA Comm. (Bansal et al., 2024) Arxiv’24 Collaboration Delegation One-by-One Synchronous Decentralized Conversation
BPMN (Ait et al., 2024) Arxiv’24 Link Collaboration Coordination One-by-One Asynchronous Decentralized Conversation
Co-STORM (Jiang et al., 2024) EMNLP’24 Link Collaboration Coordination One-by-One Synchronous Decentralized Conversation
HRC Manufa. (Lim et al., 2024) IEEE’24 Collaboration Delegation One-by-One Synchronous Decentralized Conversation
A2C (Tariq et al., 2025) Arxiv’24 Link Collaboration Coordination One-by-One Asynchronous Decentralized Conversation
MindAgent (Gong et al., 2023) NAACL’24 Link Collaboration Coordination One-by-One Synchronous Centralized Conversation
Ask Before Plan (Zhang et al., 2024c) EMNLP’24 Link Collaboration Coordination One-by-One Synchronous Centralized Conversation
SOTOPIA (Zhou et al., 2024) ICLR’24 Collab./Comp./Coopetition Coordination One-by-One Synchronous Centralized Conversation
@@
DigiRL (Bai et al., 2024) NeurIPS’24 Link Collaboration Delegation One-by-One Asynchronous Decentralized Conversation
WebLINX (Lù et al., 2024) Arxiv’24 Link Collaboration Delegation One-by-One Synchronous Decentralized Conversation
Auto Agent (Pan et al., 2024b) COLM’24 Link Collaboration Delegation One-by-One Asynchronous Decentralized Conversation
WebCanvas (Pan et al., 2024d) Arxiv’24 Link Collaboration Delegation One-by-One Synchronous Decentralized Conversation
MineWorld (Guo et al., 2025) Arxiv’25 Link Collaboration Delegation One-by-One Synchronous Decentralized Conversation

### v2

@@ 7 Conclusion
This paper presents a comprehensive review of LLM-based Human-Agent Systems. We introduce a structured taxonomy covering five core dimensions: environment and profiling, human feedback, interaction types, orchestration paradigms, and communication, and use it to classify and analyze existing research on LLM-HAS. We also summarize representative implementation frameworks, benchmark datasets, and evaluation metrics to support reproducibility and comparative analysis. Finally, we identify key challenges and unresolved issues in current LLM-HAS research. These issues remain major obstacles to the development of effective, adaptive, safe and trustworthy human-agent systems. We hope this review offers a comprehensive understanding of the LLM-HAS landscape and serves as a practical guide for future research.
@@ Appendix B Human Feedback Type and Subtype
In this appendix, we present a detailed overview of human feedback types and their subtypes, as summarized in Table B.1. This table provides concise definitions and illustrates how humans provide feedback to LLM-based agents in LLM-HAS. While the main paper introduced the broad categories of evaluative, corrective, guidance, and implicit feedback, here we expand each category into more granular subtypes, ranging from scalar ratings and preference rankings to direct edits, demonstrations, and inferred behavioral signals. Recognizing these subtypes clarifies the ways in which humans interact with LLM agents, by offering precise instructions and well-defined tasks, to enhance the accuracy and quality of generated outputs. This deeper understanding empowers users to optimize their interactions with LLM-based agents. Additionally, the systematic breakdown of human feedback provides a foundation for cross-study comparisons. It underscores the diverse strategies through which human users can guide, correct, or collaborate with LLM-based agents in a more detailed way.
@@
Human Action / Control Human directly takes actions and control. Collaborates with humans to effectively finish tasks or learns from human actions.
Table B.1: Human Feedback Types and Subtypes. The subtypes of evaluative feedback includes preference ranking, scalar rating, and binary assessment. The subtypes of corrective feedback includes the direct edits or refinement. The subtypes of guidance feedback includes the demonstration and instructions or critiques. The subtypes of implicit feedback include the human action or control.
@@ Appendix C Evaluation Metrics
In this section, we introduce evaluation metrics specifically designed for human-agent systems across four key aspects: feedback mechanisms, adaptability, trust and safety, and communication methods. To evaluate feedback mechanisms, Liu et al. (2024b) assesses a human-robot teaming framework using multi-modal language feedback at varying frequency levels (inactive, passive, active, superactive). Metz et al. (2024) proposes seven metrics, expressiveness, ease, definiteness, context independence, precision, unbiasedness, and informativeness, to evaluate feedback quality. In the education domain, Seßler et al. (2025) adopts six dimensions based on educational feedback theory. Spencer et al. (2020) evaluates the Expert Intervention Learning (EIL) method by comparing robot performance with and without expert intervention. For adaptability, Hauptman et al. (2023) examines how human-LLM agents respond to cyber incidents under different levels of autonomy across five NIST-defined phases. For trust and safety, Levy et al. (2024) introduces a benchmark that evaluates web agents on their ability to comply with policies, avoid unsafe behavior, respect security constraints, and handle errors gracefully, including seeking user input when needed. Finally, Karten et al. (2023) assess four categories of communication methods in human-agent teaming, focusing on effectiveness and interpretability within simulated environments of Predator-PreyLowe et al. (2017) and Traffic JunctionSingh et al. (2018).
In addition to these aspects, AXIS Lu et al. (2024) and SYNERGAI Chen et al. (2024b) evaluate the effectiveness and robustness of human–LLM agent systems in the domains of operating systems and embedded AI, respectively. These studies highlight how evaluation criteria can vary significantly depending on the specific task or application context, reflecting differences in system constraints, performance expectations, and interaction complexity.
@@ Appendix D Tables
Table D.1: ① Environment Configuration and ② Human Feedback to LLM-based agents in human–agent systems. Environment Configuration specifies whether a single or multiple humans collaborate with one or more LLM-based agents, while Human Feedback characterizes the type, subtype, granularity, and interaction phase of the human feedback to the LLM-based agents.
@@
Collaborative Gym (Shao et al., 2024) Arxiv’24 Link Single Single Corrective, Guidance Refinements, Instructions Segment During Task
@@
MTOM (Zhang et al., 2024b) Arxiv’24 Single Single Implicit Human Action Segment During Task
@@
FineArena (Xu et al., 2025) Arxiv’25 Single Multiple Guidance Demonstrations Segment, Holistic Initial Setup, During Task
@@
Prison Dilemm (Jiang et al., 2025) Arxiv’25 Single Single Implicit Human Action Segment During Task
@@
InteractGen (Sun et al., 2024b) THU’24 Multiple Multiple Guidance Demonstrations Segment Initial Setup, During Task
AI Chains (Wu et al., 2022) CHI’24 Single Single Corrective Refinements Segment During Task
Drive As You Speak (Cui et al., 2024) WACV’24 Single Single Guidance Demonstrations Holistic Initial Setup
@@
AgentCoord (Pan et al., 2024a) Arxiv’24 Link Single Multiple Evaluative, Corrective Preference Ranking, Refinements Segment, Holistic Initial Setup, During Task
@@
CowPilot (Huq et al., 2025) Arxiv’25 Link Single Single Corrective, Evaluative Binary Assessment, Refinements Segment During Task
@@
EasyLAN (Pan et al., 2024b) Arxiv’24 Single Multiple Corrective, Guidance Demonstrations, Refinements Segment, Holistic During Task
@@
Hierarchical Agent (Liu et al., 2023b) AAMAS’24 Single Multiple Guidance Demonstrations Segment During Task
@@
SWEET-RL (Zhou et al., 2025) Arxiv’25 Link Single Single Corrective, Implicit Refinements, Human Action Segment Initial Setup, During Task
@@
HRC Assembly (Gkournelos et al., 2024) CIRP’24 Single Multiple Guidance Demonstrations Segment During Task
@@
REVECA (Seo et al., 2025) Arxiv’24 Single Multiple Guidance Demonstrations Holistic Initial Setup
@@
AssistantX (Sun et al., 2024a) Arxiv’24 Link Multiple Multiple Implicit, Guidance Human Action, Demonstrations Holistic, Segment Initial Setup, During Task
@@
MINT (Wang et al., 2024b) ICLR’24 Link Multiple Single Evaluative, Corrective, Guidance Binary Assessment, Refinements, Instructions Holistic During Task
Help Feedback (Mehta et al., 2024) EACL’24 Single Single Evaluative, Guidance Demonstrations, Instructions, Binary Assessment Holistic, Segment During Task
@@
ConvCodeWorld (Han et al., 2025) ICLR’25 Link Single Single Guidance, Evaluative Demonstrations, Instructions, Binary Assessment Segment, Holistic During Task
ReHAC (Feng et al., 2024) ACL’24 Link Single Single Corrective Refinements Segment During Task
@@
DPT Agent (Zhang et al., 2025) Arxiv’25 Link Single Single Guidance Instructions Holistic During Task
HRC Manipulation (Liu et al., 2023a) IEEE’23 Single Single Corrective, Guidance Demonstrations, Refinements Segment During Task
HRC DMP (Liu et al., 2024a) IEEE’24 Single Single Corrective, Guidance Refinements, Demonstrations Segment During Task
PARTNR (Chang et al., 2024) ICLR’25 Link Single Single Guidance Demonstrations Holistic Inital Setup
Organized Teams (Guo et al., 2024b) Arxiv’24 Link Single Multiple Guidance Demonstrations Holistic, Segment Initial Setup, During Task
CoELA (Zhang et al., 2024a) ICLR’23 Single Multiple Guidance Demonstrations Holistic, Segment Initial Setup, During Task
Agency Task (Sharma et al., 2024) EACL’24 Link Single Single Guidance Demonstrations Segment During Task
GDfC (Wang et al., 2025c) SME’25 Single Multiple Guidance, Evaluative Demonstrations, Binary Assessment, Preference Ranking Holistic, Segment Initial Setup, During Task, Post Task
PDFChatAnnotator (Tang et al., 2024) IUI’24 Single Single Corrective, Guidance Demonstrations, Refinements Segment During Task
Attentive Supp. (Tanneberg et al., 2024a) IEEE’24 Link Multiple Single Implicit, Guidance Demonstrations, Human Action Segment During Task
HRC Trust (Ye et al., 2023) IEEE’23 Single Single Guidance Demonstrations, Instructions Segment During Task
BPMN (Ait et al., 2024) Arxiv’24 Link Multiple Multiple Guidance, Corrective Instructions, Refinements Segment During Task, Post Task
Co-STORM (Jiang et al., 2024) EMNLP’24 Link Single Multiple Guidance Demonstrations Segment During Task
HRC Manufa. (Lim et al., 2024) IEEE’24 Single Single Corrective, Guidance Demonstrations, Refinements, Instructions Segment Initial Setup, During Task
A2C (Tariq et al., 2025) Arxiv’24 Link Multiple Multiple Guidance, Evaluative Binary Assessment, Instructions Holistic, Segment During Task
MindAgent (Gong et al., 2023) NAACL’24 Link Single Multiple Guidance Demonstrations Segment During Task
Ask Before Plan (Zhang et al., 2024c) EMNLP’24 Link Single Multiple Guidance Demonstrations Segment Initial Setup, During Task
SOTOPIA (Zhou et al., 2024) ICLR’24 Multiple Multiple Evaluative, Implicit Scaler Rating, Human Action Holistic, Segment During Task, Post Task
PaLM-(Driess et al., 2023) ICML’23 Link Single Single Guidance, Implicit Demonstrations, Human Action Holistic, Segment Initial Setup, During Task
TaPA (Wu et al., 2023) Arxiv’23 Link Single Single Guidance Demonstrations Holistic, Segment Initial Setup
MetaGPT (Hong et al., 2023) ICLR’24 Link Single Multiple Guidance Demonstrations Holistic Initial Setup
@@
DigiRL (Bai et al., 2024) NeurIPS’24 Link Single Single Evaluative, Guidance Binary Assessment, Demonstrations Holistic During Task, Post Task
WebLINX (Lù et al., 2024) Arxiv’24 Link Single Multiple Guidance Demonstrations Holistic, Segment Initial Setup, During Task
MineWorld (Guo et al., 2025) Arxiv’25 Link Multiple Single Implicit Human Action Segment During Task
M3HF (Wang et al., 2025d) ICML’25 Multiple Multiple Evaluative, Guidance Binary Assessment, Instructions Segment, Holistic During Task, Post Task
SMALL (Wang et al., 2024c) Arxiv’24 Single Multiple Guidance Instructions Segment Initial Setup
MAIH (Wang et al., 2024c) Arxiv’24 Single Single Implicit Human Action Segment During Task, Post Task
HRT-ML (Liu et al., 2024b) Arxiv’24 Single Multiple Corrective, Guidance, Implicit Refinements, Instructions, Human Action Holistic Initial Setup, During Task, Post Task
AXIS (Lu et al., 2024) Arxiv’25 Single Multiple Evaluative, Implicit, Corrective Human Action, Refinements, Binary Assessment Holistic Initial Setup, During Task, Post Task
EmoAgent (Qiu et al., 2025) Arxiv’25 Single Multiple Corrective, Implicit, Guidance Human Action, Instructions, Binary Assessment Segment, Holistic During Task, Post Task
SymbioticRAG (Sun et al., 2025) Arxiv’25 Single Single Corrective, Implicit, Evaluative Binary Assessment, Refinements, Demonstrations, Instructions, Human Action Segment Initial Setup, During Task, Post Task
Table D.2: ① Interaction ② Orchestration ③ Communication in LLM-based human–agent systems. Interaction types capture the human and agent collaboration type; Orchestration covers task strategy and temporal synchronization; Communication describes how messages are structured and delivered in the system.
@@
Paper Venue Code/ Data Types Variant Strategy Synchronization Structure Mode
Collaborative Gym (Shao et al., 2024) Arxiv’24 Link Collaboration Cooperation, Delegation One-by-One Asynchronous Decentralized Conversation
MTOM (Zhang et al., 2024b) Arxiv’24 Collaboration Coordination, Cooperation Simultaneous Synchronous Decentralized Conversation
FineArena (Xu et al., 2025) Arxiv’25 Collaboration Delegation, Cooperation One-by-One Synchronous Hierarchical Conversation
Prison Dilemm (Jiang et al., 2025) Arxiv’25 Coopetition One-by-One Asynchronous Decentralized Conversation
@@
AI Chains (Wu et al., 2022) CHI’24 Collaboration Cooperation One-by-One Synchronous Decentralized Conversation
Drive As You Speak (Cui et al., 2024) WACV’24 Collaboration Delegation One-by-One Synchronous Centralized Conversation
AgentCoord (Pan et al., 2024a) Arxiv’24 Link Collaboration Coordination One-by-One Synchronous Hierarchical Conversation
@@
EasyLAN (Pan et al., 2024b) Arxiv’24 Collaboration Delegation, Supervision One-by-One Synchronous Hierarchical Observation
Hierarchical Agent (Liu et al., 2023b) AAMAS’24 Collaboration Supervision, Delegation, Cooperation One-by-One Synchronous Hierarchical Conversation
SWEET-RL (Zhou et al., 2025) Arxiv’25 Link Collaboration Delegation One-by-One Synchronous Centralized Conversation
HRC Assembly (Gkournelos et al., 2024) CIRP’24 Collaboration Delegation, Cooperation One-by-One Synchronous Decentralized Conversation
REVECA (Seo et al., 2025) Arxiv’24 Collaboration Cooperation One-by-One Synchronous Decentralized Conversation
AssistantX (Sun et al., 2024a) Arxiv’24 Link Collaboration Delegation, Cooperation One-by-One Asynchronous Decentralized Message Pool
MINT (Wang et al., 2024b) ICLR’24 Link Collaboration Delegation, Cooperation One-by-One Synchronous Decentralized Conversation
Help Feedback (Mehta et al., 2024) EACL’24 Collaboration Supervision, Delegation, Cooperation One-by-One Asynchronous Decentralized Conversation
@@
ReHAC (Feng et al., 2024) ACL’24 Link Collaboration Coordination, Supervision One-by-One Synchronous Decentralized Conversation
DPT Agent (Zhang et al., 2025) Arxiv’25 Link Collaboration Coordination Simultaneous Asynchronous Decentralized Observation
@@
HRC DMP (Liu et al., 2024a) IEEE’24 Collaboration Delegation, Supervision One-by-One Synchronous Decentralized Conversation
PARTNR (Chang et al., 2024) ICLR’25 Link Collaboration Coordination, Cooperation Simultaneous Synchronous Decentralized, Centralized Observation
Organized Teams (Guo et al., 2024b) Arxiv’24 Link Collaboration Cooperation, Coordination One-by-One Synchronous Decentralized, Centralized, Hierarchical Conversation
CoELA (Zhang et al., 2024a) ICLR’23 Collaboration Cooperation, Coordination Simultaneous Synchronous Decentralized Conversation
@@
BPMN (Ait et al., 2024) Arxiv’24 Link Collaboration Coordination Simultaneous Asynchronous Decentralized Message Pool
Co-STORM (Jiang et al., 2024) EMNLP’24 Link Collaboration Coordination One-by-One Synchronous Centralized Conversation
HRC Manufa. (Lim et al., 2024) IEEE’24 Collaboration Delegation, Cooperation One-by-One Synchronous Centralized Conversation
A2C (Tariq et al., 2025) Arxiv’24 Link Collaboration Cooperation One-by-One Asynchronous Hierarchical Conversation
MindAgent (Gong et al., 2023) NAACL’24 Link Collaboration Coordination Simultaneous Synchronous Centralized Conversation
Ask Before Plan (Zhang et al., 2024c) EMNLP’24 Link Collaboration Coordination, Delegation One-by-One Synchronous Hierarchical Conversation
SOTOPIA (Zhou et al., 2024) ICLR’24 Collaboration, Competition, Coopetition Coordination, Cooperation One-by-One Synchronous Decentralized Conversation
@@
DigiRL (Bai et al., 2024) NeurIPS’24 Link Collaboration Delegation One-by-One Synchronous Centralized Conversation
WebLINX (Lù et al., 2024) Arxiv’24 Link Collaboration Delegation One-by-One Synchronous Hierarchical Conversation
MineWorld (Guo et al., 2025) Arxiv’25 Link Collaboration Delegation One-by-One Synchronous Decentralized Observation
M3HF (Wang et al., 2025d) ICML’25 Collaboration Cooperation One-by-One, Simultaneous Synchronous Centralized Message Pool
SMALL (Wang et al., 2024c) Arxiv’24 Collaboration Delegation One-by-One Asynchronous Hierarchical Message Pool
MAIH (Wang et al., 2024c) Arxiv’24 Collaboration Delegation, Cooperation, Coordination One-by-One, Simultaneous Asynchronous Decentralized, Hierarchical Message Pool
HRT-ML (Liu et al., 2024b) Arxiv’24 Collaboration Coordination, Cooperation One-by-One, Simultaneous Asynchronous, Synchronous Hierarchical, Centralized Message Pool, Conversation
AXIS (Lu et al., 2024) Arxiv’25 Collaboration Delegation One-by-One Synchronous Centralized Conversation, Observation
EmoAgent (Qiu et al., 2025) Arxiv’25 Collaboration Supervision, Coordination, Cooperation One-by-One Synchronous Hierarchical, Centralized Conversation, Observation
SymbioticRAG (Sun et al., 2025) Arxiv’25 Collaboration Cooperation, Supervision, Delegation One-by-One Synchronous Centralized Conversation